Raysun Inc, Sells micro solar panels for $1.99 each. The marketing department prepared the following semiannual forecast:

January 30,000 panels

February . 38,000 panels

March 39,000 panels

April . 41,000 panels

May 43,000 panels

June 49,000 panels

Over the past few years Raysun Inc. has collected 25% of sales in the month of sale. 60 % in the month following the month of sale 14% in the second month following month of sale.

Required:

A. Prepare a monthly sale budgets for January through June. Includes a total column.

B. Prepare a monthly cash receipts budget for the second quarter (April, May and June). Include a total column.
Answer #1


Answer 1 Monthly sales budget

Raysun Inc
Sales Budget
For the first Half from January through June
Sales Budget January February March April May June Total
Sales In Units 30000 38000 39000 41000 43000 49000 240000
Sale on Account @ 1.99 $        59,700 $        75,620 $        77,610 $          81,590 $        85,570 $        97,510 $        477,600

Answer 2 Cash Receipts budget for April, May and June

Raysun Inc
Monthly Cash Receipts Budget
April May June Total
From February Sale $ 10,586.80 0 0 $    10,586.80
From March Sale $    46,566.0 $ 10,865.40 0 $    57,431.40
From April Sale $ 20,397.50 $    48,954.0 $ 11,422.60 $    80,774.10
From May Sale 0 $ 21,392.50 $    51,342.0 $    72,734.50
From June Sale 0 0 $ 24,377.50 $    24,377.50
Total Collections from Customers $ 77,550.30 81211.9 87142.1 $ 245,904.30

All above Calculations are based on followings

Collections in April
From February Sale =75620*14% $ 10,586.80
From March Sale =77610*60% $ 46,566.00
From April Sale =81590*25% $ 20,397.50
Collection in May
From March Sale =77610*14% $ 10,865.40
From April Sale =81590*60% $ 48,954.00
From May Sale =85570*25% $ 21,392.50
Collection in June
From April Sale =81590*14% $ 11,422.60
From May Sale =85570*60% $ 51,342.00
From June Sale =97510*25% $ 24,377.50

Hit Thumbs up if satisfied

Any query mention in Comment Sections please

Thank you